# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ 7 Layout Design Optimization
The layout design problem aims at seeking for a balance between facility costs and long-term operating costs. A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs. Considering transportation costs and supply stability, we analyze the following scenario: The warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$. The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$. The facility costs $C_{f}$ and operations costs $C_{o}$ are calculated as follows.
@@
Parameter $P_{s}$ and $P_{l}$ in equation (30) denote the discounted site rental cost per square meter and equipment cost per loading station, while $P_{w}$ and $P_{r}$ in equation (31) denote the discounted unit labor cost and unit rental cost of robot. The objective of the problem is modeled as follows:
$C_{d}(T^{\sigma},\theta^{\sigma},N_{o})=C_{f}+C_{o}$ (32)
where parameter $N_{o}$ denotes the minimum number of outlets that should be covered. The trade-off between these two types of costs in the objective function is the focus of the site planning stage. We formulate the layout design problem (LDP) as following integer programming model.
@@
$n_{w}^{\sigma}\leq n_{l}\qquad\forall\sigma\in\mathcal{S}$ (36)
$n_{h}=2k_{h}\ ,\quad k_{h}\in\mathcal{Z}_{+}$ (37)
$n_{v}=2k_{v}\ ,\quad k_{v}\in\mathcal{Z}_{+}$ (38)
$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+}^{2\left|\mathcal{S}\right|}$ (39)
Constraint (33) ensures that the system throughput meets predicted sorting requirements in different periods. Constraint (34) restricts the lower bound of the site size to accommodate an adequate number of outlets. Constraint (35) represents the maximum number of loading stations accommodated in the aisle network under RC-S. Constraint (36) limits the maximum number of workers to be stationed. Constraints (37)-(38) ensure the number of aisles even. The remaining constraint (39) requires the variables to be positive integers.
It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems. In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP-relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. To ensure the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, shown as follows.
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}$ (40)
At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up. The details of PSLP algorithm are shown in appendix C.
@@ 8 Numerical Examples for Layout Design
In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$, and labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes. Due to increased competition in the robot market, prices remain relatively stable across different companies. Consequently, we do not investigate variations in robot rental cost during the experiments, nor do we examine equipment cost at loading stations. We consider a 5-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and then the discounted costs of different components in LDP are calculated by:
@@
where $M_{i}$ denotes the unit cost of component $i$ in a month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. Labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from website for job posts (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from Geekplus company.
We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$. The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week-long shopping event. The default values of parameters is shown in table 8 and the configurations of RC-S keeps the same with table 2. Our algorithms are coded in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
Table 8: Default values of parameters in section 8
@@
8.1 Sensitivity Analysis on the site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of site floor space per square meter varies from $10\ CNY$ to $30\ CNY$. The results are presented in Figure 19(a) and Figure 19(b). The findings reveal that:
@@ 8.1 Sensitivity Analysis on the site rental cost
Figure 19: Comparison of total costs and facility costs proportion on different site rental cost
Table 9: Results of scenario $M_{s}=10,N_{o}=100$
@@
(×$(\times{10}^{6})$ (×$(\times{10}^{6})$ (×$(\times{10}^{6})$ site rental cost (%)
@@
  • When the target throughput level is low, the total costs in scenario 1 is significantly lower than that in scenario 2. This is because, to accommodate more outlets, the latter requires renting a larger initial space. In both scenarios, the curve of the optimal system cost exhibits a turning point $T^{*}$. When $T^{H}>T^{*}$, the cost escalation rate increase significantly. The turning point indicates that the initial space in the current scenario is no longer sufficient to accommodate a larger throughput, necessitating an expansion of the aisle network scale. In scenario 1, this turning point occurs at $T^{*}=9,000$; whereas in scenario 2, the turning point is at $T^{*}=21,000$.
  • The proportion of facility costs increases as $M_{s}$ rises, particularly remarkable at lower sorting throughput levels. As the throughput level becomes higher, this proportion gradually decreases and stabilizes after the turning point $T^{*}$ in each case. By observing the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to the increase in the site rental cost. Table 9 shows the optimal system configuration and corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$. When $T^{H}>T^{*}$, the optimal solutions show not only assigning more workers and robots to loading stations but also simultaneously expanding the size of the network to alleviate traffic pressure.
  • Comparing the data for $T^{H}=15,000$ and $T^{H}=30,000$ in Table 9, the total costs for the latter are nearly three times those for the former. This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the total costs growth rate of the RSS system exceeds the throughput growth rate.
From the results of study in site rental cost, the derived managerial insights are summarized as follows.
@@
The advantage of RSS is evident in its low initial investment in scenarios with both throughput demand and sortation category low.
@@
High density lead to decreased efficiency of workers. Instead of running RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
@@
Figure 20: Comparison of total costs and labor costs proportion on different unit labor cost
8.2 Sensitivity Analysis on the labor cost
In a similar manner as our previous experiments, we now explore the influence of different labor costs on the total costs in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost of each worker varies from $5,000\ CNY$ to $12,000\ CNY$. The proportions of labor costs and robot rental costs are both visualized in Figure 20(a) and Figure 20(b). Our observations are concluded as follows:
  • As $T^{H}$ increases, the proportion of labor costs gradually rises and eventually stabilizes. In every scenario, the proportion of labor costs in the optimal solution for $T^{H}=30,000$ exceeds 50%.
@@ 8.2 Sensitivity Analysis on the labor cost
From the result of study in labor cost, we summarize the following managerial insight:
@@
At lower levels of throughput, facility costs are the primary expense of RSS, whereas at higher levels of throughput, labor costs becomes the primary expense.
Additionally, the cost structure undergoes notably changes with varying sorting demands. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to tackle the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, necessitate substantial initial investments but experience only slight increments in total costs as sorting demands grow (Zou et al.,, 2021). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit price of cost components (Russell and Meller,, 2003), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
@@
Due to the traffic issues of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
@@ 9 Conclusion
This study conducts a comprehensive analysis of the RSS in modern warehouses, encompassing the underlying robot traffic management, efficiency analysis, as well as the cost composition and layout design of the system. By incorporating the throughput estimation, this research sheds light on the prediction of performance and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in logistics industry.
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovative RC-S scheme, which serves as a framework for managing a considerable number of robots simultaneously and efficiently. We provide a detailed description of the composition of phases and the method for setting parameters in RC-S. Additionally, we present a mathematical programming model FPA to minimize the travel costs at each cycle, and a heuristic algorithm as online solver. Subsequently, a theoretical analysis of the efficiency of RC-S is conducted, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with classical cooperative A* algorithm. The simulation results indicate that our control method achieves a higher level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. Results show that queueing models can experience significant distortion in some scenarios, while our model fits the real traffic flows of robots more precisely.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between the initial investment of facility $C_{f}$, and the discounted operations costs $C_{o}$. We analyze the cost composition of the RSS, and propose a layout design optimization model LDP that minimizes the total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions under distinct throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total costs. However, at higher throughput, the expenditure on labor emerges as the predominant cost element. Results also validate the discussions on the properties of our model. The key insights can guide managers in understanding the investment and returns of applying the RSS, thereby reducing the total costs of the warehouse.
For future work, there are two directions to explore. Firstly, optimizing layout design considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness. Secondly, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
@@
$\mathcal{C}$ set of all cycles
$\mathcal{R}$ set of all feasible path
$\mathcal{R}(i,j,k)$ set of feasible path connecting loading station $i$ and $j$, passing outlet $k$
$\mathcal{R}_{i},k$ set of feasible path starting at loading station $i$ and passing by outlet $k$
^$\hat{\mathcal{R}}_{i}$ set of feasible path ending at loading station $i$
@@
${T}^{\sigma}$ target throughput level in operating period $\sigma$
$\theta^{\sigma}$ Ratio of operating period $\sigma$
@@
$c_{i,k}^{r}$ travel time of feasible path $r\in R_{i,k}$
$\hat{c}_{i,k}$ penalty of delay of a sorting task from loading station $i$ to outlet $k$
$\delta_{i,k}^{r,\nu,l}$ incidence between feasible path $r\in R_{i,k}$, node $\nu$ and cycle $l$
$N_{\nu}^{l}$ remaining capacity in node $\nu$ in cycle $l$
$d_{i,k}$ sorting demand from loading station $i$ to outlet $k$
@@
$x_{i,k}^{r}$ decision on whether feasible path $r\in R_{i,k}$ is reserved in current cycle
$\hat{x}_{i,k}$ decision on whether the robot located at loading station $i$ should wait until the next cycle, with the target outlet $k$
@@
$n_{w}^{\sigma}$ number of workers in operating period $\sigma$
$n_{r}^{\sigma}$ number of robots in operating period $\sigma$

### v4

@@ 7 Layout Design Optimization
The layout design problem aims to balance facility costs and long-term operating costs. A larger facility will achieve higher maximum throughput; however, it will also increase construction and equipment costs. Considering transportation costs and supply stability, we analyze the following scenario: A warehouse signs a long-term RaaS contract with a robotics company, with each year divided into several operating periods denoted by $\sigma$. The proportion of each period’s duration within the year is represented by $\theta^{\sigma}$, and the planned throughput level during each period is $T^{\sigma}$. The facility cost $C_{f}$ and expected operations cost $C_{o}$ are calculated as follows:
@@
Parameters $P_{s}$ and $P_{l}$ in Equation (31) denote the discounted site rental cost per square meter and equipment cost per loading station, respectively, while $P_{w}$ and $P_{r}$ in equation (32) denote the discounted unit labor cost and the unit rental cost of a robot. The objective of the problem is modeled as follows:
$C_{d}=C_{f}+C_{o}$ (33)
The trade-off between these two types of costs in the objective function is the focus of the site planning stage. We formulate the layout design problem (LDP) as the following integer programming model:
@@
$\tilde{T}_{O}(n_{h},n_{v},n_{w}^{\sigma},n_{r}^{\sigma})\geq T^{\sigma},\qquad\forall\sigma\in\mathcal{S}$ (34)
@@
$n_{w}^{\sigma}\leq n_{l},\qquad\forall\sigma\in\mathcal{S}$ (37)
$n_{h}=2k_{h}\ ,\qquad k_{h}\in\mathcal{Z}_{+}$ (38)
$n_{v}=2k_{v}\ ,\qquad k_{v}\in\mathcal{Z}_{+}$ (39)
$n_{w}^{\sigma},n_{r}^{\sigma}\in\mathcal{Z}_{+},\qquad\forall\sigma\in\mathcal{S}$ (40)
Constraint (34) ensures that the system throughput meets the sorting requirements in different periods. Constraint (35) imposes a lower bound on the site size to accommodate an sufficient number of outlets $N_{o}$. Constraint (36) represents the maximum number of loading stations that can be accommodated in the aisle network under RC-S. Constraint (37) limits the maximum number of workers to be stationed. Constraints (38)-(39) ensure that the numbers of aisles are even. The remaining constraint (40) requires the decision variables to be positive integers.
It is difficult to solve LDP because of the non-linearity of the inequality constraints, especially for large-scale problems. In this research, we apply the method of penalty successive linear programming (PSLP) to solve the LP relaxation of LDP, which exhibits good robustness and convergence properties for large-scale problems (bazaraa2013nonlinear). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem augmented with the penalty function, and utilizes the concept of a trust region (updated at each iteration) to control the step size. To ensure that the constraints continuous and differentiable, we introduce additional constraints requiring that the number of robots is less than the number of available VPs, given by
$n_{r}^{\sigma}-\beta(n_{h},n_{v})\cdot(1-{(1-\frac{n_{w}^{\sigma}}{n_{h}+n_{v}})}^{2})\cdot\frac{2\tau_{e}\cdot[n_{h}(n_{v}-1)+n_{v}(n_{h}-1)]}{\tau_{c}}\leq 0\quad\forall\sigma\in\mathcal{S}.$ (41)
At the end of the algorithm, the optimal solution of the relaxed problem is converted into a feasible solution for the original problem by rounding up. The details of the PSLP algorithm are shown in Appendix C.
@@ 8 Numerical Examples for Layout Design
In this section, we conduct a sensitivity analysis on the site rental cost per square meter $P_{s}$ and the labor cost per man-month $P_{w}$, focusing on the trend of total costs under different scenarios as the sorting throughput level changes. Due to increased competition in the robot market, prices remain relatively stable across different companies. Consequently, we do not investigate variations in the robot rental cost in the experiments, nor do we examine the equipment cost at loading stations. We consider a five-year investment plan with a monthly interest rate $\gamma_{0}=0.5\%$, and the discounted costs of different components in LDP are calculated as:
@@
where $M_{i}$ denotes the unit cost of component $i$ per month. We estimate the monthly warehousing rental cost using JD Logistics’ financial reports for the first three quarters of 2023. The labor cost is sourced from data related to warehousing job recruitment in Beijing, obtained from a job-posting website (58.com). The data regarding the equipment cost of loading stations and the rental cost of robots are obtained from the Geekplus company.
We assume that the warehouse only expands its sorting capacity during a few major shopping events, meaning that there are two typical levels of sorting demand, namely the average off-peak season demand ${T}^{L}$ and the average peak season demand ${T}^{H}$, with ${T}^{L}=0.8\cdot{T}^{H}$. The ratios of the two periods are set as $(\theta^{L},\theta^{H})=(5/6,1/6)$, which means that each quarter typically includes a two-week shopping event. The default values of the parameters are shown in Table 9 and the configurations of RC-S remain the same as in Table 3. Our algorithms are implemented in Python 3.11 with the solver Gurobi Optimizer 9.0, and all experiments are conducted on a personal computer running Windows 11 with an Intel i7-9700F CPU and 16 GB of RAM. All processes are run single-threaded.
Table 9: Default parameter values in Section 8
@@
8.1 Sensitivity analysis on site rental cost
We begin by investigating the impact of site rental cost under two different scenarios: one with $N_{o}=100$ (scenario 1) and another with $N_{o}=400$ (scenario 2). The monthly rental cost of floor space per square meter varies from 10 CNY to 30 CNY. The results are presented in Figures 20(a) and  20(b). The findings reveal that:
@@ 8.1 Sensitivity analysis on site rental cost
Figure 20: Comparison of total costs and facility cost proportions on different site rental costs
Table 10: Results for scenario $M_{s}=10$, $N_{o}=100$
@@
$(\times{10}^{6})$ $(\times{10}^{6})$ $(\times{10}^{6})$ site rental cost (%)
@@
  • When the target throughput level is low, the total cost in scenario 1 is significantly lower than that in scenario 2. This is because, to accommodate more outlets, the latter requires renting a larger initial space. In both scenarios, the curve of the optimal system cost exhibits a turning point $T^{*}$. When $T^{H}>T^{*}$, the cost escalation rate increases significantly. The turning point indicates that the initial space in the current scenario is no longer sufficient to accommodate a larger throughput, necessitating an expansion of the aisle network scale. In scenario 1, this turning point occurs at $T^{*}=9{,}000$, whereas in scenario 2, the turning point is at $T^{*}=21{,}000$.
  • The proportion of facility costs increases as $M_{s}$ increases, and this effect is particularly pronounced at lower sorting throughput levels. As the throughput level becomes higher, this proportion gradually decreases and stabilizes after the turning point $T^{*}$ in each case. By examining the system configurations in each dataset, we find that this phenomenon arises because, at higher throughput demands, efficiency improvements necessitate expanding the space, which leads to an increase in the site rental cost. Table 10 shows the optimal system configuration and the corresponding costs in the case of $M_{s}=10$ and $N_{o}=100$. When $T^{H}>T^{*}$, the optimal solutions not only assign more workers and robots to the loading stations, but also expand the size of the network to alleviate traffic pressure.
  • Comparing the data for $T^{H}=15{,}000$ and $T^{H}=30{,}000$ in Table 10, the total cost for the latter is nearly three times that for the former. This indicates that due to the simultaneous growth of $C_{f}$ and $C_{o}$, the growth rate of the total cost exceeds the growth rate of throughput.
From the results of the site rental cost analysis, the derived managerial insights are summarized as follows.
@@
The advantage of RSS is evident in its low initial investment in scenarios where both throughput demand and the number of sortation categories are low.
@@
High density leads to decreased worker efficiency. Instead of running an RSS at full capacity, it is more cost-effective to appropriately expand the site size and reduce the proportion of activated loading stations.
@@
Figure 21: Proportions of labor cost and robot rental cost under different unit labor costs
8.2 Sensitivity analysis on labor cost
Similarly to our previous experiments, we now explore the influence of different labor costs on the total cost in two scenarios with $N_{o}=100$ and $N_{o}=400$. The monthly labor cost per worker varies from 5,000 CNY to 12,000 CNY. The proportions of labor cost and robot rental cost are both visualized in Figures 21(a) and 21(b). Our observations are summarized as follows:
  • As $T^{H}$ increases, the proportion of labor cost gradually increases and eventually stabilizes. In every scenario, the proportion of labor costs in the optimal solution for $T^{H}=30{,}000$ exceeds 50%.
@@ 8.2 Sensitivity analysis on labor cost
From the labor cost analysis, we summarize the following managerial insight:
@@
At lower levels of throughput, facility cost is the primary expense of RSS, whereas at higher levels of throughput, labor cost becomes the primary expense.
Additionally, the cost structure undergoes notable changes with varying sorting demand. In scenarios with high sorting demand, the proportion and growth rate of robot rental cost become more significant. This reveals a distinctive feature of RSS systems: Robots, as carriers of parcels, become less efficient in larger sites because they have to cover increased distances. Merely enhancing management methods is insufficient to address the challenge of diminishing traffic efficiency. Today’s commonly used conveyor sorters, equipped with high-power motors, require substantial initial investments but experience only slight increases in total cost as the sorting demand grows (zou2021robotic). Furthermore, the optimal system configurations of a conveyor sorter are not easily affected by fluctuations in the unit prices of cost components (russell2003cost), which makes it suitable for long-term operation in large-scale sorting scenarios. Therefore, we conclude the following insight regarding the application of RSS:
@@
Due to the traffic limitations of robots, RSS is suitable for application in small-scale scenarios, such as distribution centers at the end of the supply chain.
@@ 9 Conclusion
This study conducts a comprehensive analysis of RSS in modern warehouses, encompassing robot traffic management, system efficiency, and the cost composition and layout design of the system. By incorporating throughput estimation, this research provides guidance for performance prediction and resource allocation in modern warehouses or factories that utilize numerous robots, promoting the ongoing development of automation in the logistics industry.
In the operations stage, efficient robot scheduling emerges as a crucial element for unlocking the business value of RSS. We propose an innovative RC-S scheme, which serves as a framework for managing a large number of robots simultaneously and efficiently. We provide a detailed description of the phase structure and the parameter-setting method in RC-S. Additionally, we present a mathematical programming model, FPA, to minimize the travel cost at each cycle, and a heuristic algorithm as an online solver. We then conduct a theoretical analysis of the efficiency of RC-S, exploring the impact of various system configurations on throughput. In the validation section, we first compare our proposed traffic management framework with the classical cooperative A* algorithm. The simulation results indicate that our control method achieves a higher level of performance and computational efficiency. Furthermore, we validate the throughput estimation formula. The results show that queueing models can produce significant errors in some scenarios, whereas our model fits the real traffic flows of robots more precisely.
In the site planning stage, warehouse managers need to make decisions that involve a trade-off between initial facility investment $C_{f}$ and discounted operations costs $C_{o}$. We analyze the cost composition of RSS and propose a layout design optimization model, LDP, that minimizes total system costs. Through a detailed sensitivity analysis examining the cost dynamics of site rental and labor, we investigate cost proportions at different throughput levels. Specifically, at lower throughput, the rental cost for the site represents a considerable proportion of the total cost. However, at higher throughput, labor expenditure becomes the predominant cost component. The results also support our analysis of the model properties. Key insights can guide managers in understanding the investment and returns associated with the application of RSS, thus reducing the total cost of the warehouse.
For future work, there are two directions to explore. First, optimizing layout design by considering non-uniform demand and various outlet distributions would further improve the model’s effectiveness. Second, extending the analysis to cover the entire process from order fulfillment to sorting in modern warehouses would provide valuable insights for enhancing system performance.
@@
$\mathcal{C}$ set of all cycles starting from the current cycle
$\mathcal{R}$ set of all feasible paths
$\mathcal{R}(i,j,k)$ set of feasible paths connecting loading station $i\in\mathcal{L}$ and $j\in\mathcal{L}$, passing outlet $k\in\mathcal{O}$
$\mathcal{R}_{i,k}$ set of feasible path patterns starting from loading station $i\in\mathcal{L}$ and passing by outlet $k\in\mathcal{O}$, independent of starting cycle
$\mathcal{R}_{i}^{(0)}$ Set of feasible paths available to loading station $i$ at the beginning of current cycle, based on the current destination outlet of the pending sorting task. If no task is pending, $\mathcal{R}_{i}^{(0)}=\emptyset$
$\hat{\mathcal{R}}_{i}$ set of feasible paths ending at loading station $i\in\mathcal{L}$
@@
${T}^{\sigma}$ target throughput level in operating period $\sigma\in\mathcal{S}$
$\theta^{\sigma}$ Ratio of operating period $\sigma\in\mathcal{S}$
@@
$c_{i}^{r}$ completion time of feasible path $r\in\mathcal{R}_{i}^{(0)}$, measured in number of cycles
$\hat{c}_{i}$ penalty for delaying the assignment of a feasible path to a sorting task from loading station $i\in\mathcal{L}$ until the next cycle
$\delta_{i}^{r,t,\nu,l}$ incidence between feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$, node $\nu\in\mathcal{V}$ and cycle $l\in\mathcal{C}$
$N_{\nu}^{l}$ remaining capacity in node $\nu\in\mathcal{V}$ in cycle $l\in\mathcal{C}$
$d_{i}$ sorting demand from loading station $i\in\mathcal{L}$, $d_{i}\in\{0,1\}$
@@
$x_{i}^{r,t}$ decision on whether feasible path $r\in\mathcal{R}_{i}^{(0)}$ starting at cycle $t\in\mathcal{C}$ is reserved
$\hat{x}_{i}$ decision on whether the sorting task from loading station $i\in\mathcal{L}$ should be postponed to the next cycle
@@
$n_{w}^{\sigma}$ number of workers in operating period $\sigma\in\mathcal{S}$
$n_{r}^{\sigma}$ number of robots in operating period $\sigma\in\mathcal{S}$